OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -2 and passes through (5,-1)? A. y + 1 = -2(x-5) B. y - 1= -2(x-5) C. y - 5 = -2(x +1 ) D. y-5 = -2(x-1)

OpenStudy (campbell_st):

using the point slope formula \[y - y _{1}= m(x - x _{1})\] \[y _{1} = -1......x _{1} = 5\] and m = -2 just substitute to find the equation